OpenStudy (anonymous):

~NEED ASAP! The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under 3 years free 3 to 12 years $32 Over 12 years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

the answer choices are: A. (54 + x)32 + (82 - (54 + x)) ⋅ 44 = C B.(54 - x)32 + (82 - 54) ⋅ 44 = C C.(82 + x)32 + (82 - (54 + x)) ⋅ 44 = C D.(82 - x)32 + (82 - 54) ⋅ 44 = C

OpenStudy (anonymous):

Cost of under 3s + Cost of 3-12s + Cost of Over12s = C What's the Cost of under 3s?

OpenStudy (anonymous):

its free

OpenStudy (anonymous):

right. so we can forget about that. it's just 0. What about the cost of 3-12s. That will equal the number of 3-12 times the cost per 3-12. The cost from the table is 32. What's an expression for the number of 3-12s...given that we know there are 54 kids 12 and under and x kids are under 3?

OpenStudy (anonymous):

54 kids are 12 and under...but we don't want the under 3s..so we need to minus them out of the 54...so...the number of 3-12s is just: 54 - x